### v1

@@ Acknowledgements
We thank Peter Hase and Archiki Prasad for useful feedback and suggestions regarding experiments. This work was supported by NSF-CAREER Award 1846185, NSF-AI Engage Institute DRL-2112635, DARPA MCS Grant N66001-19-2-4031, Accelerate Foundation Models Research program, and a Google PhD Fellowship. The views contained in this article are those of the authors and not of the funding agency.
Broader Impacts
@@ Appendix A Implementation Details of MAGDi
MAGDi-distilled models are fine-tuned using Low-Rank Adaptation (LoRA) for efficiency Hu et al. (2022). We set the rank to 16 and alpha to 32. We fine-tune the student model for 10 epochs using a learning rate of 5e-6 and a batch size of 16. For the hyperparameters $\alpha$, $\beta$ and $\gamma$, please refer to our code implementation for detailed settings in each dataset. All of our experiments are run on four RTX A6000 with 48G memory each.

### v2

@@ Acknowledgements
We thank Peter Hase, Archiki Prasad, and the anonymous reviewers for useful feedback and suggestions regarding experiments. This work was supported by NSF-CAREER Award 1846185, NSF-AI Engage Institute DRL-2112635, DARPA MCS Grant N66001-19-2-4031, Accelerate Foundation Models Research program, and a Google PhD Fellowship. The views contained in this article are those of the authors and not of the funding agency.
Impact Statement
@@ Appendix A Implementation Details of MAGDi
MAGDi-distilled models are fine-tuned using Low-Rank Adaptation (LoRA) for efficiency (Hu et al., 2022). We set the rank to 16 and alpha to 32. We fine-tune the student model for 10 epochs using a learning rate of 5e-6 and a batch size of 16. For the hyperparameters $\alpha$, $\beta$ and $\gamma$, please refer to our code implementation for detailed settings in each dataset. All of our experiments are run on four RTX A6000 with 48G memory each.
@@ Appendix B Further Analyses of MAGDi
MAG edges effectively model the discussion structure.   Recall that we defined directed edges in MAGs according to the interaction pattern between agents across rounds. To further demonstrate the utility of these edges, we compare our directed MAGs (D-MAG) to fully-connected MAGs (FC-MAG) where every node is connected to every other node – and undirected MAGs (UD-MAG) where the edges follow the interaction but are undirected. As shown in Table LABEL:tab:graph_ablation, FC-MAG performs significantly worse, showing the utility of defining edges according to interaction patterns. Finally, the directionality of the edges has marginal impact on the final performance.
@@
Table 10: Dataset-wise breakdown of the performance and efficiency trade-off shown in Fig 4.
StrategyQA CSQA ARC-c GSM8K MATH Avg
# Tok (ReConcile) 924.5 936.9 448.3 642.3 1900.1 970.4
# Tok (MAGDi) 107.5 104.2 86.4 141.6 645.0 216.9
Reduction 8.6x 9.0x 5.2x 4.5x 2.9x 4.5x
Acc (ReConcile) 79.0 74.7 93.5 85.3 41.0 74.7
Acc (MAGDi) 74.2 72.6 72.6 52.3 12.8 56.9
Table 11: Utility of defining edges according to interaction structure in MAGs. Fully-connected MAGs perform worse than defining (directed or undirected) edges based on interactions.
Method StrategyQA CSQA ARC-c GSM8K MATH
FC-MAG 71.59 71.22 69.38 50.29 11.02
UD-MAG 74.02 72.61 72.46 52.10 12.66
D-MAG 74.24 72.56 72.61 52.27 12.76